# Are All Marine Species Created Equal? Performance Disparities in Underwater Object Detection

### v1

@@ 4.5 Results: Bounding Box Placement
Holothurian $0.7115$ 0.11-0.11 $0.5042$ 0.15-0.15
Echinus $0.7478$ 0.07-0.07 $0.5967$ 0.05-0.05
Scallop $0.4415$ 0.38-0.38 $0.3499$ 0.30-0.30
Starfish $0.8213$ $0.00$ $0.6507$ $0.00$
Table 3: The per-class mAP results of the balanced dataset and the deviation from the best‐performing class. The scallop class under-performance is less pronounced with higher IoU thresholds.
Specifically, when requiring better bounding box accuracy via a higher IoU threshold, the performance discrepancy becomes less between classes. This implies that localization precision is similar across classes, and the main cause of difference at lower thresholds is due to detection difficulty, not box alignment. This is underlined in Fig. 4: the balanced datasets have rarely any localization errors, i.e. if the object is not missed, the predicted bounding box usually is accurate enough. While there are some minor differences in the precision of bounding boxes across classes (Fig. 5), with the smaller and visually less distinct scallop objects having slightly larger center offsets and higher area errors more frequently, all classes follow a very similar error distribution regardless of imbalanced or balanced setup.
@@
Figure 5: Accuracy of predicted bounding boxes in balanced setup. (a) Bounding box center offset, with minor deviations but overall similar distribution across classes and datasets. (b) The bounding box area error, which highlights larger fluctuations for the small scallop objects but a similar trend overall.
@@ 5 Classification Study
This section describes our analysis on classification stage of object detection. Sec. 5.1 describes the methodology, before we explain the datasets (Sec. 5.2) and implementation (Sec. 5.3). Sec. 5.4 then highlights what factors impact how detected targets are assigned to the correct target class.
@@
Classification DUO Reference Set 0.952 0.991 0.943 0.982 0.970 0.994 0.916 0.964 0.961 0.993 0.929 0.973
Imbalanced Minimum 0.884 0.978 0.776 0.959 0.897 0.988 0.703 0.921 0.890 0.983 0.733 0.940
Balanced Full 0.842 0.990 0.664 0.955 0.952 0.969 0.962 0.929 0.893 0.979 0.786 0.942
Classification DUO Reference Set + Weighted Sampling 0.950 0.993 0.946 0.987 0.979 0.995 0.934 0.966 0.964 0.994 0.940 0.976
Table 4: Classification precision, recall and F1 score (refer to Sec. 3 for definitions) across different dataset distributions. Results highlight a precision-recall tradeoff: balanced datasets improve scallop recall, imbalanced datasets result in higher precision. Weighted sampling mitigates scallop under-performance best.
@@ 5.1 Method
In the classification experiments, we focus purely on how well the query targets are assigned to the correct class. To isolate this stage from localization effects, we extract single-class image crops that simulate the output of a bounding box regressor and feed them directly to a classifier. This setup enables controlled manipulation of the training samples to establish the impact of data quantity and class imbalance.
@@ 5.2 Datasets
Given that DUO was designed for object detection, we modified the data for our classification experiments to obtain single-object images. First, we cropped square images of every object with at least 10px padding around the bounding box. We randomly selected 20% of the training data as the validation set and removed all cropped images where the original bounding box overlapped by more than 10% with any other annotated object, thereby avoiding crops that contain multiple species. We call this cleaned subset the classification DUO reference set, which maintains the original imbalanced class distribution (69% echinus, 20% starfish, 9% holothurian, 2% scallop).
This serves as the basis for further subsets: (i) balanced sets with randomly down-sampled classes so that all provide the same data quantity as scallops, referred to as Balanced Full; (ii) reduced sets in which a) the total data quantity from the classification DUO reference set is decreased for an imbalanced minimum set of the same size as the balanced set maintaining class distribution, referred to as Imbalanced Minimum and b) the total data quantity from the balanced set is gradually decreased keeping equal instance count per class; and (iii) differently distributed sets where only samples of one class at a time are gradually reduced. We also conduct an experiment on the classification DUO reference set with weighted sampling to address the class imbalance.
@@ 5.3 Implementation
We used a ResNet-18 [12] pre-trained on ImageNet as our classifier using PyTorch. We fine-tune this model for 30 epochs with a batch size of 32. All images are resized to 224 x 224 pixels and normalized using ImageNet means and standard deviation. The training images are augmented with random horizontal flip and random rotation within $\pm$15 degrees. We use standard cross-entropy loss and Adam as optimizer with an initial learning rate of 0.001 that we reduce by the factor 0.5 every time no validation improvement has been recorded for 3 epochs, stopping at the minimum learning rate of 0.00003. All classification experiments were conducted on an Apple MacBook Pro M4 with MPS.
@@ 5.4 Results
Tab. 4 reveals a fundamental precision-recall tradeoff based on the training distribution. For scallops, imbalanced training yields higher precision but lower recall, while balanced training increases recall by 4.6% (from 91.6% to 96.2%) but significantly reduces precision by 27.9% from 94.3% to 66.4%. This pattern is most pronounced for scallops, with other classes showing greater robustness across training strategies. Scallop under-performance persists across all sets but is almost negligible when weighted sampling is applied. When the data is quantitatively balanced (Balanced Full), all classes achieve comparable and high recall but significant performance disparities are evident in precision: Fig. 6 reveals that even with balanced training, scallops achieve 55-70% precision versus >>95% for echinus and starfish, demonstrating intrinsic visual challenges.
This demonstrates that while there are benefits for larger training datasets, classification performance differences caused by inherent data characteristics will persist. The scallop class exhibits both the lowest overall precision, and the largest performance fluctuations as data is reduced, suggesting that scallop features are inherently less distinctive and more sensitive to data availability.
Figure 6: Per class classification performance on incrementally reduced balanced subsets (Solid lines: Precision; dashed lines: Recall). Recall is consistently high across all classes but there are significant gaps in precision, with the scallop class experiencing the largest fluctuations with reduced data.
Fig. 7 shows that reducing training data for other classes significantly impacts scallop classification performance, even when scallop training data remains constant. Interestingly, reducing starfish data to 30% causes an 8.2% drop in scallop precision and a 1.1% reduction in scallop recall, while reducing holothurian data results in a 7.9% precision decrease for scallops.
@@
Figure 7: Effect of data reductions on performance of scallop class. From the classification reference set (100%), training samples are reduced to 30% one class at a time. Scallop precision declines as training data for other classes (particularly starfish and holothurian) is reduced, despite keeping scallop data constant. The model relies on sufficient negative examples to understand class boundaries.
@@ 6 Discussion and Conclusion
Our key findings include: (1) Performance gaps persist under balanced training conditions, indicating intrinsic rather than data-driven challenges; (2) Class interdependence effects show that models learn target identity partly through negative examples from other species; (3) Training distribution creates fundamental precision-recall tradeoffs that should be aligned with operational requirements; and (4) Background discrimination is the primary bottleneck in underwater detection tasks.
These insights lead to practical recommendations: applications prioritizing comprehensive detection with high recall, such as conservation monitoring, should use balanced training. Commercial operations with high precision requirements should prefer imbalanced training. As separating targets from the background represents the primary technical challenge, research efforts should prioritize improving localization over classification accuracy via targeted algorithmic improvements or the use of extra sensing modalities.
By providing this systematic analysis and releasing our evaluation code and processed datasets, we aim to foster more targeted research approaches that address the specific challenges of underwater object detection rather than applying generic computer vision solutions.

### v2

@@ 4.5 Results: Bounding Box Placement
Holothurian 0.710.71/$0.80$ 0.11-0.11/0.08-0.08 0.500.50/$0.51$ 0.15-0.15/0.06-0.06
Echinus 0.750.75/$0.82$ 0.07-0.07/0.06-0.06 0.600.60/$0.49$ 0.05-0.05/0.08-0.08
Starfish 0.820.82/$0.88$ 0.000.00/$0.00$ 0.650.65/$0.57$ 0.000.00/$0.00$
Scallop 0.440.44/$0.73$ 0.38-0.38/0.15-0.15 0.350.35/$0.49$ 0.30-0.30/0.08-0.08
Table 3: DUO / RUOD-4C per-class mAP results on the balanced dataset and the deviation from the best‐performing class. The scallop class under-performance is less pronounced with higher IoU thresholds.
Specifically, when requiring better bounding box accuracy via a higher IoU threshold, the performance discrepancy becomes smaller between classes. This implies that localization precision is similar across classes, and the main cause of difference at lower thresholds is due to detection difficulty, not box alignment. This is underlined in Fig. 4: the balanced datasets have rarely any localization errors, i.e. if the object is not missed, the predicted bounding box usually is accurate enough. While there are some minor differences in the precision of bounding boxes across classes (Fig. 5), with the smaller and visually less distinct scallop objects having slightly larger center offsets and higher area errors more frequently, all classes follow a very similar error distribution regardless of imbalanced or balanced setup.
@@
Figure 5: Accuracy of predicted bounding boxes for DUO. (a) Bounding box center offset, with minor deviations but overall similar distribution across classes and datasets. (b) The bounding box area error, which highlights larger fluctuations for the small scallop objects but a similar trend overall.
@@ 5 Classification Study
This section describes our analysis on the classification stage of object detection. Sec. 5.1 describes the methodology, we explain the datasets in Sec. 5.2, and implementation details are in Sec. 5.3. Sec. 5.4 highlights our findings on what factors impact the classification of detected targets.
@@
Cls Reference Set 95.2/95.4 99.1 /98.1 94.3 /98.2 98.2 / 97.8 97.0 / 96.7 99.4 /98.3 91.6 / 97.7 96.4 /97.0 96.1 /96.0 99.3 /98.2 92.9 /98.0 97.3 /97.4
Imbalanced Minimum 88.4 / 95.3 97.8 / 97.9 77.6 / 97.7 95.9 / 97.3 89.7 / 96.1 98.8 / 98.2 70.3 / 97.6 92.1 / 96.4 89.0 / 95.7 98.3 / 98.0 73.3 / 97.7 94.0 / 96.8
Balanced Full 84.2 / 93.7 99.0 / 97.9 66.4 / 98.1 95.5 /97.9 95.2 /97.1 96.9 / 97.5 96.2/ 97.6 92.9 / 96.3 89.3 / 95.3 97.9 / 97.8 78.6 / 97.7 94.2 / 97.1
Weighted Sampling 95.0 / 95.3 99.3/ 97.9 94.6/ 97.8 98.7/97.9 97.9/ 96.5 99.5/98.3 93.4 /97.8 96.6/ 96.5 96.4/ 95.9 99.4/ 98.1 94.0/ 97.8 97.6/ 97.2
Table 4: DUO / RUOD-4C classification precision, recall and F1 score (refer to Sec. 3 for definitions) across different dataset distributions. Results on DUO highlight a precision-recall tradeoff: balanced datasets improve scallop recall, imbalanced datasets result in higher precision. Weighted sampling mitigates scallop under-performance best. For RUOD-4C, metrics are consistently high across classes and data distributions, reflecting the near-balanced original class distribution.
@@ 5.1 Method
In the classification experiments, we focus purely on how well the query targets are assigned to the correct class. To isolate this stage from localization effects, we extract single-object image crops that simulate the output of a bounding box regressor and feed them directly to a classifier. This setup enables controlled manipulation of the training samples to establish the impact of data quantity and class imbalance.
@@ 5.2 Datasets
Given that DUO and RUOD-4C were designed for object detection, we modified the data for our classification experiments to obtain single-object images. First, we cropped square images of every object with at least 10px padding around the bounding box. We randomly selected 20% of the training data for validation and removed all cropped images where the original bounding box overlapped by more than 10% with any other annotated object, thereby avoiding crops that contain multiple species. We call this cleaned subset the classification reference set, which maintains a similar class distribution to the original data in DUO (Sec. 3), i.e.  there is a pronounced class imbalance. In RUOD-4C, the holothurian class loses proportionally a greater quantity of images, making holothurian the smallest class in this scenario.
The classification reference sets serve as the basis for further subsets: (i) balanced sets with randomly down-sampled classes so that all provide the same data quantity as the minority class, referred to as Balanced Full; (ii) reduced sets in which a) the total data quantity from the classification DUO reference set is decreased for an imbalanced minimum set of the same size as the balanced set maintaining class distribution, referred to as Imbalanced Minimum and b) the total data quantity from the balanced set is gradually decreased keeping equal instance count per class; and (iii) differently distributed sets where only samples of one class at a time are gradually reduced. We also conduct an experiment on the classification reference sets with weighted sampling to address class imbalance.
@@ 5.3 Implementation
We used a ResNet-18 [14] pre-trained on ImageNet as our classifier implemented through PyTorch. The same experiments were repeated with MobileNetV2 and VisionTransformer-B/16 as classifiers and confirmed the results obtained with ResNet-18 (for further details, please refer to the Supp. Mat.).
@@ 5.4 Results
Tab. 4 demonstrates very good classification performance in general. The only major disparities are found in the scallop precision values for DUO. In RUOD-4C, all metrics are generally high and very similar across classes and across data versions. Considering the original RUOD-4C class distribution is nearly balanced (Sec. 3), it is not surprising that imbalanced vs. balanced, full vs. minimal or unweighted vs. weighted data show no significant differences. However the pronounced class imbalance in DUO requires in-depth analysis: examination of the DUO performance reveals a fundamental precision-recall tradeoff based on the training distribution. For scallops, imbalanced training yields higher precision but lower recall, while balanced training increases recall by 4.6% (from 91.6% to 96.2%) but significantly reduces precision by 27.9% from 94.3% to 66.4%. This pattern is most pronounced for scallops, with other classes showing greater robustness across training strategies. The performance gaps in DUO become almost negligible when weighted sampling is applied. When the data is quantitatively balanced (Balanced Full), all classes achieve comparable and high recall but significant performance disparities are evident in precision: Fig. 6 reveals that even with balanced training, scallops achieve 55-70% precision, vs. >>95% for echinus and starfish, demonstrating intrinsic visual challenges.
This indicates that while there are benefits for larger training datasets, classification performance differences caused by inherent data characteristics persist. The scallop class exhibits both the lowest overall precision, and the largest performance fluctuations as data is reduced, suggesting that scallop features are inherently less distinctive and more sensitive to data availability.
Figure 6: Per class classification performance on incrementally reduced balanced DUO subsets (Solid lines: Precision; dashed lines: Recall). Recall is consistently high across all classes but there are significant gaps in precision, with the scallop class experiencing the largest fluctuations with reduced data.
Fig. 7 shows that reducing training data for other classes in DUO significantly impacts scallop classification performance, even when scallop training data remains constant. Interestingly, reducing starfish data to 30% causes an 8.2% drop in scallop precision and a 1.1% reduction in scallop recall, while reducing holothurian data results in a 7.9% precision decrease for scallops.
@@
Figure 7: Effect of data reductions on scallop performance in DUO. From the classification reference set (100%), training samples are reduced to 30% one class at a time. Scallop precision declines as training data for other classes (particularly starfish and holothurian) is reduced, despite keeping scallop data constant. The model relies on sufficient negative examples to understand class boundaries.
@@ 6 Discussion and Conclusion
Our key findings include: (1) Performance gaps persist under balanced training conditions, indicating intrinsic rather than data-driven challenges; (2) Class interdependence effects show that models learn target identity partly through negative examples from other species; (3) For imbalanced data, the training distribution creates fundamental precision-recall tradeoffs that should be aligned with operational requirements; and (4) Background discrimination is the primary bottleneck in underwater detection.
These insights lead to practical recommendations: applications prioritizing recall, such as conservation monitoring, should use balanced training, while commercial operations needing high precision should prefer imbalanced training. Since background separation is the main challenge, research should focus on improving localization over classification, using algorithmic advances or additional sensing. This systematic analysis aims to foster targeted solutions for underwater object detection.